OpenStudy (anonymous):

Integrate xe^x . Can somebody give me a step by step solution. Ans is e^x(x-1)

OpenStudy (anonymous):

Use integration by parts.\[\int u dv = uv - \int v du\] with \(u=x\) and \(dv = e^x dx\). what must \(du\) and \(v\) be? By the formula, what must this integral equal?

OpenStudy (anonymous):

\[e^x(x+1)\]

OpenStudy (anonymous):

Thanks for the hint :)

Directrix (directrix):

See attachment for tips. http://tinyurl.com/aubrttj